OpenStudy (anonymous):

the linear equation for m=3/4 and (6,4)

OpenStudy (smurfy14):

y=3x/4-1/2

OpenStudy (anonymous):

what is the formula.

OpenStudy (smurfy14):

point slope formula: y-y1=m(x-x1)

OpenStudy (smurfy14):

when you plug in #s it is y-4=3/4(x-6)